# Event Camera Meets Resource-Aware Mobile Computing:

### v3

@@ 6.3. Simultaneous Localization and Mapping (SLAM)
Event-based SLAM leverages the high temporal resolution and low latency of event cameras for robust localization and mapping in dynamic environments [259, 260]. A key enhancement is multimodal fusion, combining complementary sensors to overcome individual limitations. For instance, Ultimate SLAM [261] fuses events, frames, and IMU data for HDR and high-speed scenarios, while Implicit Event-RGBD Neural SLAM [260] integrates event and RGB-D data to handle motion blur and lighting changes. Advances in feature representation further boost robustness. Line-based SLAM methods [158] mitigate feature loss in fast motion or low-light using PTAM frameworks, while optimization-based approaches like CMax-SLAM [262] apply contrast maximization for precise rotational motion estimation through event-based global bundle adjustment. Despite these promising results, most event-based SLAM methods remain at the proof-of-concept stage, validated mainly in controlled or semi-controlled scenarios. Only a few methods have been demonstrated in real-world dynamic environments [259, 263], indicating that full deployment readiness is still limited.
@@ 7. Future direction and discussion
Despite extensive research on event cameras, their application in mobile sensing remains in its early stages, with significant opportunities for advancement while balancing trade-offs in latency, power consumption, and accuracy.
@@
(1) Improving event cameras using optics devices. Event cameras respond only to illumination changes and remain inactive in fully static scenes, limiting continuous perception. While some hardware-based solutions exist, they often reduce energy efficiency due to added mechanical components [35, 264]. Future work could explore dynamic optical elements, such as electro-optic materials, to induce controlled illumination changes via optical phase arrays, enabling detection in static environments. Hybrid optical-electronic systems, combining event cameras with active illumination or computational imaging, may further enhance performance [265, 249]. Key challenges remain, including designing diverse scanning patterns without blind spots, developing easily implementable non-mechanical illumination devices, and improving event signal quality under high-speed illumination.
@@
(2) Designing neuromorphic hardware to facilitate event processing. Current hardware faces challenges in processing event-based data. CPUs suffer from frequent context switching, while GPUs are ill-suited for asynchronous, high-frequency events. FPGAs provide parallelism and low latency but lack end-to-end pipeline optimization. Although some efforts have developed dedicated accelerators [174, 39, 171], the massive data volume in event stream (e.g., thousands of events in milliseconds) can easily overwhelm I/O bandwidth, and end-to-end latency often violates worst-case budgets under event bursts, causing jitter that disrupts control loops. Moreover, mapping high-level neural models to neuromorphic ISAs remains manual and brittle, limiting programmability and debugging. Future work should pursue specialized neuromorphic hardware tailored for event-driven computing. Architectures based on SNNs or asynchronous pipelines can better exploit event sparsity, while higher-throughput interconnects with QoS, priority channels for control streams, and traffic shaping can alleviate I/O bottlenecks. Hard real-time scheduling, bounded-latency NoCs, and deadline-aware routers are needed to guarantee temporal determinism. Finally, stable intermediate representations (IRs) and compiler toolchains are essential to improve programmability across diverse neuromorphic platforms.
@@
(3) Leveraging the complementary strengths of event camera and other sensors. Event cameras complement frame cameras, radar, and LiDAR by providing high temporal resolution and low latency, which mitigate motion blur and enhance perception in dynamic environments. While frame cameras offer texture and color [32, 266], radar ensures robustness under poor lighting and weather [5, 267], and LiDAR supplies precise 3D geometry [268, 269], their integration with event streams enables improved visual odometry, detection, reconstruction, and SLAM. Yet, most existing fusion methods remain loosely coupled and overlook the intrinsic properties of raw multimodal data—for instance, frame images are spatially dense but temporally sparse, event and radar data are spatially sparse yet temporally dense, and LiDAR point clouds are sparse in both space and time. Future research should therefore develop tightly coupled deep learning– or optimization-based frameworks that explicitly exploit these raw data complementary characteristics, alongside optimized hardware designs for real-time, energy-efficient robotic applications.
@@
(4) Bio-inspired algorithm design. Event cameras inherently exhibit neuromorphic traits, making them well-suited for bio-inspired algorithms. Integrating SNNs, which process discrete spikes like the brain, enables sparse, low-power, and precise event-driven perception for tasks such as recognition and tracking [270]. Bio-inspired models based on primate vision improve segmentation [39], while neuromorphic control systems emulate sensorimotor loops for fast decisions [271]. Nevertheless, most current innovations primarily focus on mimicking the biological mechanisms, developing sophisticated algorithms or hardware for high-level recognition and scene understanding. While inspiring, such approaches may not be optimal for tasks requiring strict real-time performance due to their significant computational overhead, or in some cases, may be impractical to realize. Future research should therefore explore alternative nature-inspired strategies—rather than strictly replicating biological mechanisms—by designing bio-inspired models that enable efficient event processing, and by optimizing neuromorphic hardware to effectively support these systems.
@@
(5) Hardware-software co-optimization techniques. Efficient event data processing requires tight hardware-software co-design. Traditional hardware struggles with asynchronous, sparse event streams. Dedicated neuromorphic hardware—such as SNN chips, FPGA-based processors, and ASICs—can better support event-driven computation, improving efficiency [174, 39]. On the software side, event-driven programming models and lightweight architectures are needed to reduce latency and redundancy, as conventional deep learning frameworks are not well-suited for event data [32]. However, most existing approaches focus on either hardware or software in isolation, overlooking cross-layer optimizations. Techniques such as dynamic resource allocation, memory access optimization, and hardware-aware model compression could further improve speed and energy efficiency. Future research should therefore emphasize holistic co-design of algorithms and neuromorphic hardware to realize real-time, low-power event-based vision systems for robotics, autonomous driving, and edge intelligence.
@@ 8. Conclusion
Event-based vision is a transformative approach for mobile sensing, offering high temporal resolution, low latency, and energy efficiency. This survey reviews event camera principles, event representations, algorithms, hardware/software acceleration, and diverse mobile applications. Despite advantages, challenges remain in event processing, sensor fusion, and real-time use on resource-limited platforms. Future work should enhance hardware with advanced optics, develop neuromorphic processors for asynchronous data, and apply bio-inspired algorithms to boost perception. Integrating event cameras with LiDAR and radar will further broaden applications in dynamic settings. We hope this survey inspires research and practical deployment in the mobile sensing field.

### v4

@@ 6.3. Simultaneous Localization and Mapping (SLAM)
Event-based SLAM leverages the high temporal resolution and low latency of event cameras for robust localization and mapping in dynamic environments [niu2025esvo2, 10655500]. A key enhancement is multimodal fusion, combining complementary sensors to overcome individual limitations. For instance, Ultimate SLAM [8258997] fuses events, frames, and IMU data for HDR and high-speed scenarios, while Implicit Event-RGBD Neural SLAM [10655500] integrates event and RGB-D data to handle motion blur and lighting changes. Advances in feature representation further boost robustness. Line-based SLAM methods [Chamorro2022EventBasedLS] mitigate feature loss in fast motion or low-light using PTAM frameworks, while optimization-based approaches like CMax-SLAM [10474186] apply contrast maximization for precise rotational motion estimation through event-based global bundle adjustment. Despite these promising results, most event-based SLAM methods remain at the proof-of-concept stage, validated mainly in controlled or semi-controlled scenarios. Only a few methods have been demonstrated in real-world dynamic environments [niu2025esvo2, zhou2021event], indicating that full deployment readiness is still limited.
@@ 7. Future direction and discussion
Despite extensive research on event cameras, their application in mobile embodied perception remains in early stages, with significant opportunities for advancement while balancing trade-offs in latency, power consumption, and accuracy.
@@
(1) Improving event cameras using optics devices. Event cameras respond only to illumination changes and remain inactive in fully static scenes, limiting continuous perception. While some hardware-based solutions exist, they often reduce energy efficiency due to added mechanical components [he2024microsaccade, bao2024temporal]. Future work could explore dynamic optical elements, such as electro-optic materials, to induce controlled illumination changes via optical phase arrays, enabling detection in static environments. Hybrid optical-electronic systems, combining event cameras with active illumination or computational imaging, may further enhance performance [yu2024eventps, muglikar2023event]. Key challenges remain, including designing diverse scanning patterns without blind spots, developing easily implementable non-mechanical illumination devices, and improving event signal quality under high-speed illumination.
@@
(2) Designing neuromorphic hardware to facilitate event processing. Current hardware faces challenges in processing event-based data. CPUs suffer from frequent context switching, while GPUs are ill-suited for asynchronous, high-frequency events. FPGAs provide parallelism and low latency but lack end-to-end pipeline optimization. Although some efforts have developed dedicated accelerators [Cao2024EventBoost, xu2023taming, Pei2019Towards], the massive data volume in event stream (e.g., thousands of events in milliseconds) can easily overwhelm I/O bandwidth, and end-to-end latency often violates worst-case budgets under event bursts, causing jitter that disrupts control loops. Moreover, mapping high-level neural models to neuromorphic ISAs remains manual and brittle, limiting programmability and debugging. Future work should pursue specialized neuromorphic hardware tailored for event-driven computing. Architectures based on SNNs or asynchronous pipelines can better exploit event sparsity, while higher-throughput interconnects with QoS, priority channels for control streams, and traffic shaping can alleviate I/O bottlenecks. Hard real-time scheduling, bounded-latency NoCs, and deadline-aware routers are needed to guarantee temporal determinism. Finally, stable intermediate representations (IRs) and compiler toolchains are essential to improve programmability across diverse neuromorphic platforms.
@@
(3) Leveraging the complementary strengths of event camera and other sensors. Event cameras complement frame cameras, radar, and LiDAR by providing high temporal resolution and low latency, which mitigate motion blur and enhance perception in dynamic environments. While frame cameras offer texture and color [gehrig2024low, zhao2025urbanvideo], radar ensures robustness under poor lighting and weather [wang2025ultra], and LiDAR supplies precise 3D geometry [cui2022dense, zhou2024bring], their integration with event streams enables improved visual odometry, detection, reconstruction, and SLAM. Yet, most existing fusion methods remain loosely coupled and overlook the intrinsic properties of raw multimodal data—for instance, frame images are spatially dense but temporally sparse, event and radar data are spatially sparse yet temporally dense, and LiDAR point clouds are sparse in both space and time. Future research should therefore develop tightly coupled deep learning– or optimization-based frameworks that explicitly exploit these raw data complementary characteristics, alongside optimized hardware designs for real-time, energy-efficient robotic applications.
@@
(4) Bio-inspired algorithm design. Event cameras inherently exhibit neuromorphic traits, making them well-suited for bio-inspired algorithms. Integrating SNNs, which process discrete spikes like the brain, enables sparse, low-power, and precise event-driven perception for tasks such as recognition and tracking [ghosh2009spiking]. Bio-inspired models based on primate vision improve segmentation [xu2023taming], while neuromorphic control systems emulate sensorimotor loops for fast decisions [liu2022toward]. Nevertheless, most current innovations primarily focus on mimicking the biological mechanisms, developing sophisticated algorithms or hardware for high-level recognition and scene understanding. While inspiring, such approaches may not be optimal for tasks requiring strict real-time performance due to their significant computational overhead, or in some cases, may be impractical to realize. Future research should therefore explore alternative nature-inspired strategies—rather than strictly replicating biological mechanisms—by designing bio-inspired models that enable efficient event processing, and by optimizing neuromorphic hardware to effectively support these systems.
@@
(5) Hardware-software co-optimization techniques. Efficient event data processing requires tight hardware-software co-design. Traditional hardware struggles with asynchronous, sparse event streams. Dedicated neuromorphic hardware—such as SNN chips, FPGA-based processors, and ASICs—can better support event-driven computation, improving efficiency [Cao2024EventBoost, xu2023taming]. On the software side, event-driven programming models and lightweight architectures are needed to reduce latency and redundancy, as conventional deep learning frameworks are not well-suited for event data [gehrig2024low]. However, most existing approaches focus on either hardware or software in isolation, overlooking cross-layer optimizations. Techniques such as dynamic resource allocation, memory access optimization, and hardware-aware model compression could further improve speed and energy efficiency. Future research should therefore emphasize holistic co-design of algorithms and neuromorphic hardware to realize real-time, low-power event-based vision systems for robotics, autonomous driving, and edge intelligence.
@@ 8. Conclusion
Event-based vision is a transformative approach for mobile embodied perception, offering high temporal resolution, low latency, and energy efficiency. This survey reviews event camera principles, event representations, algorithms, hardware/software acceleration, and diverse mobile applications. Despite advantages, challenges remain in event processing, sensor fusion, and real-time use on resource-limited agents. Future work should enhance hardware with advanced optics, develop neuromorphic processors for asynchronous data, and apply bio-inspired algorithms to boost perception. Integrating event cameras with LiDAR and radar will further broaden applications in dynamic settings. We hope this survey inspires research and practical deployment in the mobile embodied perception field.